# Measuring Social Norms of Large Language Models

### v1

@@ Appendix D Summary of Skills
Subject: Language Arts
Skill: revise-the-sentence-using-a-stronger-verb
Description: Read the sentence.
After a difficult climb, Madelyn **went** back down the hill.
Now, revise the sentence to show that Madelyn was **clumsy**.
After a difficult climb, Madelyn ___ back down the hill.
Choices: [(a) floated, (b) stumbled, (c) plunged]
Answer index: 1
Subject: Language Arts
Skill: use-the-correct-subject-or-verb
Description: Complete the sentence with the best **verb**.
Both of us ___ the banjo, but Jenna also plays the flute.
Choices: [(a) plays, (b) play]
Answer index: 1
Subject: Language Arts
Skill: choose-between-personal-and-reflexive-pronouns
Description: Select the best pronoun to complete the sentence.
Will you please give ___ directions to the Lanberry Community Center?
Choices: [(a) me, (b) myself]
Answer index: 0
Subject: Language Arts
Skill: choose-between-subject-and-object-personal-pronouns
Description: Select the correct pronoun to complete the sentence.
When Tara is upset, she always confides in ___.
Choices: [(a) he, (b) him]
Answer index: 1
Subject: Language Arts
Skill: choose-the-best-concluding-sentence
Description: Read the paragraph. Then, add the better concluding sentence.
@@
Over the course of his career, Mel Blanc provided the voices for hundreds of cartoon characters. One of his most famous characters was Bugs Bunny. Bugs Bunny was a charming yet fearless rabbit who was famous for saying "What’s up, Doc?" Blanc described his process for how he came up with Bugs Bunny’s voice. First, he looked at sketches, or rough drawings, of the character. Then, he looked at a set of pictures called a storyboard. The storyboard showed how Bugs Bunny would behave in a cartoon. The cartoon’s creators told Blanc that Bugs Bunny’s character was a "tough little stinker." Blanc thought about all the tough-sounding accents he had heard. After some experimenting, Blanc finally brought Bugs Bunny’s voice to life. ___
Choices: [(a) Similarly, Blanc is known for creating sound effects., (b) This process clearly worked, because Bugs Bunny became one of Blanc’s most recognized characters.]
Answer index: 1
Subject: Language Arts
Skill: create-compound-sentences
Description: Form a compound sentence.
Sarah ___ at the rink.
@@
Choices: [(a) doesn’t play ice hockey, but she figure skates, (b) Montoya, famous for her graceful jumps, can be seen]
Answer index: 0
Subject: Language Arts
Skill: identify-the-narrative-point-of-view
Description: Which point of view does the narrator use in the passage?
A few days after this I had to go out with Ginger in the carriage. I wondered how we should get on together; but except laying her ears back when I was led up to her, she behaved very well.
From Anna Sewell, Black Beauty. Copyright 1999 by Dover Publications
Choices: [(a) first person, (b) second person, (c) third person]
Answer index: 0
Subject: Language Arts
Skill: is-the-word-an-adjective-or-adverb
Description: Is the word in bold an adjective or an adverb?
Zoe tried on a **violet** gown with cap sleeves.
Choices: [(a) adjective, (b) adverb]
Answer index: 0
Subject: Language Arts
Skill: is-it-a-complete-sentence-a-fragment-or-a-run-on
Description: Which is a **run-on sentence**?
@@
Choices: [(a) I mopped the floor, Mom washed all the dishes., (b) The girls and boys swung on the swings.]
Answer index: 0
Subject: Language Arts
Skill: use-actions-and-dialogue-to-understand-characters
Description: Read the sentences.
@@
Jaylen asked that the math award be given to his whole team instead of just him. "I couldn’t have done it without them," he said.
Based on these sentences, what is Jaylen probably like?
Choices: [(a) humble, (b) snobby]
Answer index: 0
Subject: Language Arts
Skill: use-adverbs-to-compare
Description: Complete the sentence with the correct form of the adverb.
Wesley can sing the ___ of the three soloists.
Choices: [(a) lower, (b) lowest]
Answer index: 1
Subject: Language Arts
Skill: which-definition-matches-the-sentence
Description: What does **bloom** mean in this sentence?
Mr. Herman said, "I think Brandon will bloom into a successful and confident young man."
Choices: [(a) **bloom** "verb" to grow, (b) **bloom** "noun" a blossom]
Answer index: 0
Subject: Language Arts
Skill: is-the-sentence-in-the-past-present-or-future-tense
Description: Which tense does the sentence use?
Erica parked her scooter in front of the store.
Choices: [(a) past tense, (b) present tense, (c) future tense]
Answer index: 0
Subject: Language Arts
Skill: formatting-and-capitalizing-titles
Description: Which sentence is correct?
@@
Choices: [(a) Almost everyone watched the TV show "Foster’s Home for Imaginary Friends.", (b) Almost everyone watched the TV show "Foster’s Home for Imaginary Friends".]
Answer index: 1
Subject: Language Arts
Skill: choose-the-antonym
Description: Which word is an antonym of **succeed**?
Choices: [(a) fail, (b) victory]
Answer index: 0
Subject: Language Arts
Skill: choose-the-best-topic-sentence
Description: Read the paragraph. Then, add the better topic sentence.
@@
___ In slacklining, people build strength and balance as they practice walking across a narrow, stretchy strip of fabric. As in tightrope walking, it takes a lot of practice to stay steady on this line. When slackliners feel confident enough, they may test their balance by slacklining higher up. Some even perform tricks on a slackline.
Choices: [(a) Some people practice yoga poses on slacklines., (b) Slacklining is a sport that uses balance and strength.]
Answer index: 1
Subject: Language Arts
Skill: read-poetry
Description: Review the poem.
My Model Car
The hours piled on,
As I sprawled out on my carpet,
Tinkering with tweezers,
Fiddling with rubber bands,
Trying to make my model car come to life.
Finally, eyes bleary, I went for a walk,
Leaving the car parts behind.
I took in the details of the neighborhood,
The trills and chirps of birds in the trees,
The thick, smoky smell of a grill
Fired up in my neighbor’s yard.
I came back and opened my door.
Hoping the car had assembled itself. No, but—
I suddenly saw the parts with new eyes,
My mind clear and ready.
My tools bent and twisted the pieces
Like they’d always known what to do.
What is the main theme of the poem?
Choices: [(a) It’s important to ask for help with a project when you need it., (b) Taking a break from a task can help you approach it in a new way., (c) Make sure you understand all the steps of your task before you start.]
Answer index: 1
Subject: Language Arts
Skill: show-character-emotions-and-traits
Description: You are writing a story about Devon, who hears some bad news. Choose the best description to show that Devon is **disappointed**.
Choices: [(a) Devon swallowed hard and tried to force herself to smile., (b) "Okay," Devon said, continuing to chop carrots.]
Answer index: 0
@@
Subject: Language Arts
Skill: identify-text-structures
Description: Read the text.
@@
Although there are many kinds of writing, most people divide texts into two different types: fiction and nonfiction. Fiction means stories that are made up. "Charlotte’s Web" is fiction because, of course, spiders and pigs don’t really talk. Nonfiction, however, gives you facts about the real world. You can find nonfiction books about history, science, real people, animals, and many other topics. Nonfiction also includes newspapers and textbooks, while fiction includes stories, plays, and novels. Reading both fiction and nonfiction can be fun and educational.
Which text structure does the text use?
Choices: [(a) cause-effect, (b) compare-contrast]
Answer index: 1
Subject: Language Arts
Skill: which-sentence-matches-the-definition
Description: Which sentence uses this meaning of **concrete**?
*concrete** "adjective" specific and real
@@
Choices: [(a) His fingerprints were **concrete** proof that he had opened the safe., (b) We can’t walk on the sidewalk because the **concrete** is still drying.]
Answer index: 0
Subject: Language Arts
Skill: determine-the-meaning-of-idioms-from-context-set-1
Description: What is the meaning of **on the tip of one’s tongue**?
I can’t quite remember the name of that magician we saw perform in Las Vegas, but it is **on the tip of my tongue**.
Choices: [(a) almost remembered, (b) a certainty]
Answer index: 0
Subject: Language Arts
Skill: use-guide-words
@@
Description: Which word would you find on a dictionary page with the following guide words?
*skate** - **squire**
Choices: [(a) spare, (b) station]
Answer index: 0
Subject: Language Arts
Skill: use-context-to-identify-the-meaning-of-a-word
Description: Read the passage and then answer the question.
@@
The Great Pyramid of Giza is the largest of the ancient Egyptian pyramids. When completed around 2560 BCE, it stood 481 feet high, as tall as a modern 44-story building. The Great Pyramid was the world’s tallest building for more than three thousand years. More than two million huge stone blocks were used to construct it. The builders of the pyramid positioned it **meticulously** so that it perfectly faces true north. Scientists aren’t sure how the ancient Egyptians knew how to place the base. Experts also disagree about how those huge blocks were moved into place. They don’t even fully understand the design of the inside of the pyramid. More work must be done to solve these mysteries.
What is the meaning of **meticulously** as used in the passage?
Choices: [(a) with many people pushing, (b) in a great hurry, (c) happily, without any hesitation, (d) carefully, with attention to detail]
Answer index: 3
Subject: Language Arts
Skill: punctuating-dialogue
Description: Which sentence is correct?
@@
Choices: [(a) "Please hire a carpenter to finish our basement", Mrs. Huffman urged., (b) "Please hire a carpenter to finish our basement," Mrs. Huffman urged.]
Answer index: 1
Subject: Language Arts
Skill: determine-the-meanings-of-similes-and-metaphors
Description: What is the meaning of the simile?
Linda’s typing was like a tap dancer performing onstage.
Choices: [(a) Linda made rapid tapping noises., (b) Linda made her performance on the stage.]
Answer index: 0
Subject: Language Arts
Skill: use-key-details-to-determine-the-main-idea
Description: Look at these details from a paragraph about the same topic:
By taking music lessons, students learn to read sheet music.
Through music lessons, students can learn about music from other countries.
Choose the main, or central, idea that ties all the details together.
Choices: [(a) Music lessons introduce students to many types of music., (b) Music lessons teach more than just how to play an instrument., (c) Piano lessons are one of the most popular types of music lesson.]
Answer index: 1
Subject: Language Arts
Skill: choose-between-adjectives-and-adverbs
Description: Choose the word that best describes the adjective in bold.
As a wedding gift, Elena gave the couple a quilt that was ___ **sewn** by hand.
Choices: [(a) full, (b) fully]
Answer index: 1
Subject: Language Arts
Skill: read-realistic-fiction
Description: Read the story.
title: The Artist’s Secret
Destiny smiled. Then she reached for her palette, so she could start on C.
Based on paragraphs 1 and 2, how does Destiny probably feel about the art project at the beginning of the story?
Choices: [(a) She cares about it turning out just right., (b) She wants to finish it as quickly as possible., (c) She views it as a competition that she is winning., (d) She is annoyed about having to mix paint colors.]
Answer index: 0
Subject: Language Arts
Skill: use-the-meanings-of-words-as-clues-to-the-meanings-of-greek-and-latin-roots
Description: A biography is a written account of someone’s life. What does the root **graph** mean?
Choices: [(a) look at or observe, (b) write or draw, (c) measure]
Answer index: 1
Subject: Language Arts
Skill: use-coordinating-conjunctions
Description: Complete the sentence with the best conjunction.
Pam wants to dye her hair purple ___ red.
Choices: [(a) or, (b) but]
Answer index: 0
Subject: Language Arts
Skill: identify-similes-and-metaphors
Description: Does the sentence use a simile or a metaphor?
Mrs. Boyer is kind, and her heart is gold.
Choices: [(a) simile, (b) metaphor]
Answer index: 1
Subject: Language Arts
Skill: use-greek-and-latin-roots-as-clues-to-the-meanings-of-words
Description: The root **vis** means "see or look at". What does the word **visual** mean?
Choices: [(a) relating to hearing or sound, (b) relating to seeing or sight, (c) relating to flavor or taste]
Answer index: 1
Subject: Language Arts
Skill: interpret-the-meaning-of-an-allusion-from-its-source
Description: Read the source of the allusion.
@@
In the fairy tale, Cinderella is a poor young woman who is treated badly. However, she later marries a prince and becomes a loved and respected princess.
What is the meaning of the allusion in the sentence below?
Choices: [(a) an amazing improvement, (b) a journey without troubles]
Answer index: 0
Subject: Language Arts
Skill: positive-and-negative-connotation
Description: Which phrase has a more **negative** connotation?
Choices: [(a) a warm room, (b) a stuffy room]
Answer index: 1
@@
Subject: Language Arts
Skill: draw-inferences-from-a-text
Description: Read the text.
Kari looked under the table. Then she searched the whole room.
"You have got to be kidding!" Kari shouted.
Grumbling, Kari tore the puzzle apart and put it back in the box. On the top of the box, she crossed out 1000 and wrote 999 above it.
Based on the text, which sentence is more likely to be true?
Choices: [(a) Kari doesn’t like building puzzles., (b) Kari’s puzzle is missing a piece.]
Answer index: 1
Subject: Language Arts
Skill: is-the-sentence-simple-or-compound
Description: Which is a **compound sentence**?
@@
Choices: [(a) Jayden grew too many pumpkins, so he sold the extras at the market., (b) British general John Burgoyne surrendered to American forces at the Battle of Saratoga.]
Answer index: 0
Subject: Language Arts
Skill: identify-elements-of-poetry
Description: Choose the poem that uses **end rhyme**.
@@
Choices: [(a) I break a staff. I break the tough branch. I know no light in the woods. I have lost pace with the winds., (b) Sign of a nation, great and strong Toward her people from foreign wrong: Pride and glory and honor,—all Live in the colors to stand or fall.]
Answer index: 1
Subject: Language Arts
Skill: read-about-business-and-technology
Description: Review the text.
title: Where Are You?
GPS has also become very useful in the field of science. Scientists have figured out ways to use GPS for monitoring earthquakes and volcanoes. GPS can also be used to measure water levels and snowfall. While working with wild animals, some animal researchers also use GPS devices. These scientists can attach small devices onto animals to track the animals’ movements and gather important data about them. Who knows how GPS will help us next?
Which is the best summary of the text?
@@
Choices: [(a) GPS receivers help people find their way. They work by sending radio signals to satellites in space. Drivers, people enjoying outdoor sports, and emergency workers all use GPS. GPS can also be used for other purposes, like for tracking animals and for science research., (b) "GPS" stands for Global Positioning System. GPS works anywhere on Earth, except deep in caves and underwater. People use GPS every day to drive and play sports. They also use GPS to find lost pets., (c) GPS devices have become very important in our world. People use them for driving and outdoor sports. Scientists use them to monitor earthquakes, volcanoes, and the phases of the moon. GPS can also be used to remind people to walk their pets.]
Answer index: 0
Subject: Language Arts
Skill: read-about-history
Description: Read the text.
title: A Whale of a Record
@@
It might seem incredible that just one album could help save the lives of so many whales. But remember that Songs of the Humpback Whale was no ordinary record. This one-of-a-kind album revealed secrets about whales that had long been hidden beneath the waves. It allowed these amazing and mysterious creatures to finally share their songs with the human world.
What is the text about?
Choices: [(a) It is about how musicians recorded an album to encourage people to ban deep-sea whaling., (b) It is about how whales learned to use sounds to communicate after biologists played them recordings., (c) It is about how people changed their mind about recordings of nature after hearing an album of whale songs., (d) It is about how an album of whale recordings helped increase protection for whales around the world.]
Answer index: 3
Subject: Language Arts
Skill: identify-plurals-singular-possessives-and-plural-possessives
Description: What is the word in bold?
Are there any **statues** in your town?
Choices: [(a) a singular possessive noun, (b) a plural noun]
Answer index: 1
Subject: Language Arts
Skill: analyze-the-effects-of-figures-of-speech-on-meaning-and-tone
Description: Read the text. The figure of speech is shown in bold.
Then I knew that even though **nobody in the world** figured that old car had any good purpose, Ob knew there was some real reason to let it sit. And when May died, he figured out what it was.
I never saw two people love each other so much.
From Cynthia Rylant, Missing May. Copyright 1992 by Cynthia Rylant
What does this example of **hyperbole** tell you?
Choices: [(a) It suggests that very few people thought the car had a purpose., (b) It suggests that the car travels around the world.]
Answer index: 0
Subject: Language Arts
Skill: determine-the-meanings-of-greek-and-latin-roots
Description: What does the root **tract** mean?
Choices: [(a) bend or curve, (b) pull or drag, (c) far away]
Answer index: 1
Subject: Language Arts
Skill: is-it-a-telling-sentence-or-an-asking-sentence
Description: What kind of sentence is this?
My father reads to me.
Choices: [(a) telling sentence, (b) asking sentence]
Answer index: 0
Subject: Language Arts
Skill: find-the-vowel-in-the-word
Description: Find the vowel.
Choices: [(a) o, (b) n]
Answer index: 0
Subject: Language Arts
@@
Skill: complete-the-sentence-with-the-correct-short-vowel-word
Description: Which word makes the most sense in the sentence?
The bug is on his ___.
Choices: [(a) nap, (b) leg]
Answer index: 1
Subject: Language Arts
Skill: complete-the-rhyme
Description: Complete the rhyme.
The men sit in the ___.
Choices: [(a) den, (b) back]
Answer index: 0
Subject: Language Arts
Skill: capitalize-the-first-letter-of-a-sentence
Description: Complete the sentence with an uppercase letter.
___ans can go fast.
Choices: [(a) V, (b) v]
Answer index: 0
Subject: Language Arts
Skill: who-what-when-where-or-why
Description: Pick the right question word.
___ did your mom go?
Choices: [(a) Who, (b) Where]
Answer index: 1
Subject: Language Arts
Skill: what-am-i
Description: Answer the riddle.
I am small.
What am I?
Choices: [(a) a baby, (b) a box]
Answer index: 0
Subject: Language Arts

### v2

@@ Appendix D Summary of Skills
Subject: Language Arts Skill: revise-the-sentence-using-a-stronger-verb Description: Read the sentence. After a difficult climb, Madelyn **went** back down the hill. Now, revise the sentence to show that Madelyn was **clumsy**. After a difficult climb, Madelyn ___ back down the hill. Choices: [(a) floated, (b) stumbled, (c) plunged] Answer index: 1
Subject: Language Arts Skill: use-the-correct-subject-or-verb Description: Complete the sentence with the best **verb**. Both of us ___ the banjo, but Jenna also plays the flute. Choices: [(a) plays, (b) play] Answer index: 1
Subject: Language Arts Skill: choose-between-personal-and-reflexive-pronouns Description: Select the best pronoun to complete the sentence. Will you please give ___ directions to the Lanberry Community Center? Choices: [(a) me, (b) myself] Answer index: 0
Subject: Language Arts Skill: choose-between-subject-and-object-personal-pronouns Description: Select the correct pronoun to complete the sentence. When Tara is upset, she always confides in ___. Choices: [(a) he, (b) him] Answer index: 1
@@
Subject: Language Arts Skill: choose-the-best-concluding-sentence Description: Read the paragraph. Then, add the better concluding sentence. Over the course of his career, Mel Blanc provided the voices for hundreds of cartoon characters. One of his most famous characters was Bugs Bunny. Bugs Bunny was a charming yet fearless rabbit who was famous for saying "What’s up, Doc?" Blanc described his process for how he came up with Bugs Bunny’s voice. First, he looked at sketches, or rough drawings, of the character. Then, he looked at a set of pictures called a storyboard. The storyboard showed how Bugs Bunny would behave in a cartoon. The cartoon’s creators told Blanc that Bugs Bunny’s character was a "tough little stinker." Blanc thought about all the tough-sounding accents he had heard. After some experimenting, Blanc finally brought Bugs Bunny’s voice to life. ___ Choices: [(a) Similarly, Blanc is known for creating sound effects., (b) This process clearly worked, because Bugs Bunny became one of Blanc’s most recognized characters.] Answer index: 1
@@
Subject: Language Arts Skill: create-compound-sentences Description: Form a compound sentence. Sarah ___ at the rink. Choices: [(a) doesn’t play ice hockey, but she figure skates, (b) Montoya, famous for her graceful jumps, can be seen] Answer index: 0
Subject: Language Arts Skill: identify-the-narrative-point-of-view Description: Which point of view does the narrator use in the passage? A few days after this I had to go out with Ginger in the carriage. I wondered how we should get on together; but except laying her ears back when I was led up to her, she behaved very well. From Anna Sewell, Black Beauty. Copyright 1999 by Dover Publications Choices: [(a) first person, (b) second person, (c) third person] Answer index: 0
Subject: Language Arts Skill: is-the-word-an-adjective-or-adverb Description: Is the word in bold an adjective or an adverb? Zoe tried on a **violet** gown with cap sleeves. Choices: [(a) adjective, (b) adverb] Answer index: 0
@@
Subject: Language Arts Skill: is-it-a-complete-sentence-a-fragment-or-a-run-on Description: Which is a **run-on sentence**? Choices: [(a) I mopped the floor, Mom washed all the dishes., (b) The girls and boys swung on the swings.] Answer index: 0
@@
Subject: Language Arts Skill: use-actions-and-dialogue-to-understand-characters Description: Read the sentences. Jaylen asked that the math award be given to his whole team instead of just him. "I couldn’t have done it without them," he said. Based on these sentences, what is Jaylen probably like? Choices: [(a) humble, (b) snobby] Answer index: 0
Subject: Language Arts Skill: use-adverbs-to-compare Description: Complete the sentence with the correct form of the adverb. Wesley can sing the ___ of the three soloists. Choices: [(a) lower, (b) lowest] Answer index: 1
Subject: Language Arts Skill: which-definition-matches-the-sentence Description: What does **bloom** mean in this sentence? Mr. Herman said, "I think Brandon will bloom into a successful and confident young man." Choices: [(a) **bloom** "verb" to grow, (b) **bloom** "noun" a blossom] Answer index: 0
Subject: Language Arts Skill: is-the-sentence-in-the-past-present-or-future-tense Description: Which tense does the sentence use? Erica parked her scooter in front of the store. Choices: [(a) past tense, (b) present tense, (c) future tense] Answer index: 0
@@
Subject: Language Arts Skill: formatting-and-capitalizing-titles Description: Which sentence is correct? Choices: [(a) Almost everyone watched the TV show "Foster’s Home for Imaginary Friends.", (b) Almost everyone watched the TV show "Foster’s Home for Imaginary Friends".] Answer index: 1
Subject: Language Arts Skill: choose-the-antonym Description: Which word is an antonym of **succeed**? Choices: [(a) fail, (b) victory] Answer index: 0
@@
Subject: Language Arts Skill: choose-the-best-topic-sentence Description: Read the paragraph. Then, add the better topic sentence. ___ In slacklining, people build strength and balance as they practice walking across a narrow, stretchy strip of fabric. As in tightrope walking, it takes a lot of practice to stay steady on this line. When slackliners feel confident enough, they may test their balance by slacklining higher up. Some even perform tricks on a slackline. Choices: [(a) Some people practice yoga poses on slacklines., (b) Slacklining is a sport that uses balance and strength.] Answer index: 1
Subject: Language Arts Skill: read-poetry Description: Review the poem. My Model Car The hours piled on, As I sprawled out on my carpet, Tinkering with tweezers, Fiddling with rubber bands, Trying to make my model car come to life. Finally, eyes bleary, I went for a walk, Leaving the car parts behind. I took in the details of the neighborhood, The trills and chirps of birds in the trees, The thick, smoky smell of a grill Fired up in my neighbor’s yard. I came back and opened my door. Hoping the car had assembled itself. No, but— I suddenly saw the parts with new eyes, My mind clear and ready. My tools bent and twisted the pieces Like they’d always known what to do. What is the main theme of the poem? Choices: [(a) It’s important to ask for help with a project when you need it., (b) Taking a break from a task can help you approach it in a new way., (c) Make sure you understand all the steps of your task before you start.] Answer index: 1
Subject: Language Arts Skill: show-character-emotions-and-traits Description: You are writing a story about Devon, who hears some bad news. Choose the best description to show that Devon is **disappointed**. Choices: [(a) Devon swallowed hard and tried to force herself to smile., (b) "Okay," Devon said, continuing to chop carrots.] Answer index: 0
@@
Subject: Language Arts Skill: identify-text-structures Description: Read the text. Although there are many kinds of writing, most people divide texts into two different types: fiction and nonfiction. Fiction means stories that are made up. "Charlotte’s Web" is fiction because, of course, spiders and pigs don’t really talk. Nonfiction, however, gives you facts about the real world. You can find nonfiction books about history, science, real people, animals, and many other topics. Nonfiction also includes newspapers and textbooks, while fiction includes stories, plays, and novels. Reading both fiction and nonfiction can be fun and educational. Which text structure does the text use? Choices: [(a) cause-effect, (b) compare-contrast] Answer index: 1
@@
Subject: Language Arts Skill: which-sentence-matches-the-definition Description: Which sentence uses this meaning of **concrete**? **concrete** "adjective" specific and real Choices: [(a) His fingerprints were **concrete** proof that he had opened the safe., (b) We can’t walk on the sidewalk because the **concrete** is still drying.] Answer index: 0
Subject: Language Arts Skill: determine-the-meaning-of-idioms-from-context-set-1 Description: What is the meaning of **on the tip of one’s tongue**? I can’t quite remember the name of that magician we saw perform in Las Vegas, but it is **on the tip of my tongue**. Choices: [(a) almost remembered, (b) a certainty] Answer index: 0
@@
Subject: Language Arts Skill: use-guide-words Description: Which word would you find on a dictionary page with the following guide words? **skate** - **squire** Choices: [(a) spare, (b) station] Answer index: 0
@@
Subject: Language Arts Skill: use-context-to-identify-the-meaning-of-a-word Description: Read the passage and then answer the question. The Great Pyramid of Giza is the largest of the ancient Egyptian pyramids. When completed around 2560 BCE, it stood 481 feet high, as tall as a modern 44-story building. The Great Pyramid was the world’s tallest building for more than three thousand years. More than two million huge stone blocks were used to construct it. The builders of the pyramid positioned it **meticulously** so that it perfectly faces true north. Scientists aren’t sure how the ancient Egyptians knew how to place the base. Experts also disagree about how those huge blocks were moved into place. They don’t even fully understand the design of the inside of the pyramid. More work must be done to solve these mysteries. What is the meaning of **meticulously** as used in the passage? Choices: [(a) with many people pushing, (b) in a great hurry, (c) happily, without any hesitation, (d) carefully, with attention to detail] Answer index: 3
@@
Subject: Language Arts Skill: punctuating-dialogue Description: Which sentence is correct? Choices: [(a) "Please hire a carpenter to finish our basement", Mrs. Huffman urged., (b) "Please hire a carpenter to finish our basement," Mrs. Huffman urged.] Answer index: 1
Subject: Language Arts Skill: determine-the-meanings-of-similes-and-metaphors Description: What is the meaning of the simile? Linda’s typing was like a tap dancer performing onstage. Choices: [(a) Linda made rapid tapping noises., (b) Linda made her performance on the stage.] Answer index: 0
Subject: Language Arts Skill: use-key-details-to-determine-the-main-idea Description: Look at these details from a paragraph about the same topic: By taking music lessons, students learn to read sheet music. Through music lessons, students can learn about music from other countries. Choose the main, or central, idea that ties all the details together. Choices: [(a) Music lessons introduce students to many types of music., (b) Music lessons teach more than just how to play an instrument., (c) Piano lessons are one of the most popular types of music lesson.] Answer index: 1
Subject: Language Arts Skill: choose-between-adjectives-and-adverbs Description: Choose the word that best describes the adjective in bold. As a wedding gift, Elena gave the couple a quilt that was ___ **sewn** by hand. Choices: [(a) full, (b) fully] Answer index: 1
Subject: Language Arts Skill: read-realistic-fiction Description: Read the story. title: The Artist’s Secret Destiny smiled. Then she reached for her palette, so she could start on C. Based on paragraphs 1 and 2, how does Destiny probably feel about the art project at the beginning of the story? Choices: [(a) She cares about it turning out just right., (b) She wants to finish it as quickly as possible., (c) She views it as a competition that she is winning., (d) She is annoyed about having to mix paint colors.] Answer index: 0
Subject: Language Arts Skill: use-the-meanings-of-words-as-clues-to-the-meanings-of-greek-and-latin-roots Description: A biography is a written account of someone’s life. What does the root **graph** mean? Choices: [(a) look at or observe, (b) write or draw, (c) measure] Answer index: 1
Subject: Language Arts Skill: use-coordinating-conjunctions Description: Complete the sentence with the best conjunction. Pam wants to dye her hair purple ___ red. Choices: [(a) or, (b) but] Answer index: 0
Subject: Language Arts Skill: identify-similes-and-metaphors Description: Does the sentence use a simile or a metaphor? Mrs. Boyer is kind, and her heart is gold. Choices: [(a) simile, (b) metaphor] Answer index: 1
Subject: Language Arts Skill: use-greek-and-latin-roots-as-clues-to-the-meanings-of-words Description: The root **vis** means "see or look at". What does the word **visual** mean? Choices: [(a) relating to hearing or sound, (b) relating to seeing or sight, (c) relating to flavor or taste] Answer index: 1
@@
Subject: Language Arts Skill: interpret-the-meaning-of-an-allusion-from-its-source Description: Read the source of the allusion. In the fairy tale, Cinderella is a poor young woman who is treated badly. However, she later marries a prince and becomes a loved and respected princess. What is the meaning of the allusion in the sentence below? Choices: [(a) an amazing improvement, (b) a journey without troubles] Answer index: 0
Subject: Language Arts Skill: positive-and-negative-connotation Description: Which phrase has a more **negative** connotation? Choices: [(a) a warm room, (b) a stuffy room] Answer index: 1
@@
Subject: Language Arts Skill: draw-inferences-from-a-text Description: Read the text. Kari looked under the table. Then she searched the whole room. "You have got to be kidding!" Kari shouted. Grumbling, Kari tore the puzzle apart and put it back in the box. On the top of the box, she crossed out 1000 and wrote 999 above it. Based on the text, which sentence is more likely to be true? Choices: [(a) Kari doesn’t like building puzzles., (b) Kari’s puzzle is missing a piece.] Answer index: 1
@@
Subject: Language Arts Skill: is-the-sentence-simple-or-compound Description: Which is a **compound sentence**? Choices: [(a) Jayden grew too many pumpkins, so he sold the extras at the market., (b) British general John Burgoyne surrendered to American forces at the Battle of Saratoga.] Answer index: 0
@@
Subject: Language Arts Skill: identify-elements-of-poetry Description: Choose the poem that uses **end rhyme**. Choices: [(a) I break a staff. I break the tough branch. I know no light in the woods. I have lost pace with the winds., (b) Sign of a nation, great and strong Toward her people from foreign wrong: Pride and glory and honor,—all Live in the colors to stand or fall.] Answer index: 1
@@
Subject: Language Arts Skill: read-about-business-and-technology Description: Review the text. title: Where Are You? GPS has also become very useful in the field of science. Scientists have figured out ways to use GPS for monitoring earthquakes and volcanoes. GPS can also be used to measure water levels and snowfall. While working with wild animals, some animal researchers also use GPS devices. These scientists can attach small devices onto animals to track the animals’ movements and gather important data about them. Who knows how GPS will help us next? Which is the best summary of the text? Choices: [(a) GPS receivers help people find their way. They work by sending radio signals to satellites in space. Drivers, people enjoying outdoor sports, and emergency workers all use GPS. GPS can also be used for other purposes, like for tracking animals and for science research., (b) "GPS" stands for Global Positioning System. GPS works anywhere on Earth, except deep in caves and underwater. People use GPS every day to drive and play sports. They also use GPS to find lost pets., (c) GPS devices have become very important in our world. People use them for driving and outdoor sports. Scientists use them to monitor earthquakes, volcanoes, and the phases of the moon. GPS can also be used to remind people to walk their pets.] Answer index: 0
@@
Subject: Language Arts Skill: read-about-history Description: Read the text. title: A Whale of a Record It might seem incredible that just one album could help save the lives of so many whales. But remember that Songs of the Humpback Whale was no ordinary record. This one-of-a-kind album revealed secrets about whales that had long been hidden beneath the waves. It allowed these amazing and mysterious creatures to finally share their songs with the human world. What is the text about? Choices: [(a) It is about how musicians recorded an album to encourage people to ban deep-sea whaling., (b) It is about how whales learned to use sounds to communicate after biologists played them recordings., (c) It is about how people changed their mind about recordings of nature after hearing an album of whale songs., (d) It is about how an album of whale recordings helped increase protection for whales around the world.] Answer index: 3
Subject: Language Arts Skill: identify-plurals-singular-possessives-and-plural-possessives Description: What is the word in bold? Are there any **statues** in your town? Choices: [(a) a singular possessive noun, (b) a plural noun] Answer index: 1
Subject: Language Arts Skill: analyze-the-effects-of-figures-of-speech-on-meaning-and-tone Description: Read the text. The figure of speech is shown in bold. Then I knew that even though **nobody in the world** figured that old car had any good purpose, Ob knew there was some real reason to let it sit. And when May died, he figured out what it was. I never saw two people love each other so much. From Cynthia Rylant, Missing May. Copyright 1992 by Cynthia Rylant What does this example of **hyperbole** tell you? Choices: [(a) It suggests that very few people thought the car had a purpose., (b) It suggests that the car travels around the world.] Answer index: 0
Subject: Language Arts Skill: determine-the-meanings-of-greek-and-latin-roots Description: What does the root **tract** mean? Choices: [(a) bend or curve, (b) pull or drag, (c) far away] Answer index: 1
Subject: Language Arts Skill: is-it-a-telling-sentence-or-an-asking-sentence Description: What kind of sentence is this? My father reads to me. Choices: [(a) telling sentence, (b) asking sentence] Answer index: 0
Subject: Language Arts Skill: find-the-vowel-in-the-word Description: Find the vowel. Choices: [(a) o, (b) n] Answer index: 0
@@
Subject: Language Arts Skill: complete-the-sentence-with-the-correct-short-vowel-word Description: Which word makes the most sense in the sentence? The bug is on his ___. Choices: [(a) nap, (b) leg] Answer index: 1
Subject: Language Arts Skill: complete-the-rhyme Description: Complete the rhyme. The men sit in the ___. Choices: [(a) den, (b) back] Answer index: 0
Subject: Language Arts Skill: capitalize-the-first-letter-of-a-sentence Description: Complete the sentence with an uppercase letter. ___ans can go fast. Choices: [(a) V, (b) v] Answer index: 0
Subject: Language Arts Skill: who-what-when-where-or-why Description: Pick the right question word. ___ did your mom go? Choices: [(a) Who, (b) Where] Answer index: 1
Subject: Language Arts Skill: what-am-i Description: Answer the riddle. I am small. What am I? Choices: [(a) a baby, (b) a box] Answer index: 0